OpenStudy (anonymous):

The area of a rectangle is 14 square meters. Find the length and width of the rectangle if its length is 5 meters greater than its width. Use an equation and the formula for the Area of a rectangle= (width)(length).

OpenStudy (anonymous):

A = w * L w * ( w + 5) = 14 w^2 + 5w - 14 = 0 Can you continue from here?

OpenStudy (anonymous):

Yes I think so! Thank you so much!

OpenStudy (anonymous):

No questions :)